One number is 3 units larger than twice another number. If their sum is zero, find the two numbers?
Answer by ankor@dixie-net.com(22740)

One number = x, and another number = y
:
One number is 3 units larger than twice another number.
x = 2y + 3
:
If their sum is zero,
x + y = 0
:
find the two numbers?
Substitute (2y+3) for x in the equation, x + y = 0
(2y+3) + y = 0
2y + y + 3 = 0
3y = -3
y = -1
:
Find x:
x + (-1) = 0
x = +1